### v1

# DvD: Unleashing a Generative Paradigm for Document Dewarping via Coordinates-based Diffusion Model

Abstract.
Document dewarping aims to rectify deformations in photographic document images, thus improving text readability, which has attracted much attention and made great progress, but it is still challenging to preserve document structures. Given recent advances in diffusion models, it is natural for us to consider their potential applicability to document dewarping. However, it is far from straightforward to adopt diffusion models in document dewarping due to their unfaithful control on highly complex document images (e.g., 2000$\times$3000 resolution). In this paper, we propose DvD, the first generative model to tackle document Dewarping via a Diffusion framework. To be specific, DvD introduces a coordinate-level denoising instead of typical pixel-level denoising, generating a mapping for deformation rectification. In addition, we further propose a time-variant condition refinement mechanism to enhance the preservation of document structures. In experiments, we find that current document dewarping benchmarks can not evaluate dewarping models comprehensively. To this end, we present AnyPhotoDoc6300, a rigorously designed large-scale document dewarping benchmark comprising 6,300 real image pairs across three distinct domains, enabling fine-grained evaluation of dewarping models. Comprehensive experiments demonstrate that our proposed DvD can achieve state-of-the-art performance with acceptable computational efficiency on multiple metrics across various benchmarks including DocUNet, DIR300, and AnyPhotoDoc6300. The new benchmark and code will be publicly available.
Photographic Documents Images, Document Unwarping, Document Dewarping, Diffusion Model, Optical Character Recognition, Generative AI
submissionid: aaaaccs: Applied computing Document analysisccs: Applied computing Optical character recognitionccs: Applied computing Document scanningccs: Computing methodologies Image processing
1. Introduction
With the ubiquity of smartphones, taking photos to digitize documents has become an increasingly convenient and common practice. However, compared with the scanned documents, document images captured in daily scenes often suffer from poor readability caused by severe geometric deformations (e.g., folds, curves, crumples, etc.). These deformations hinder both human readability and specialist optical character recognition (OCR) engines, even multimodal large language models (MLLMs) (Scius-Bertrand et al., 2024). To improve readability, document dewarping task emerges to restore flat documents before downstream document digitization pipelines (e.g., Layout analysis, Text spotting) (Li et al., 2025; Wan et al., 2024), achieving comparable OCR performance to flat counterparts.
Figure 1. Comparisons to the existing paradigm for document dewarping. Unlike the existing paradigms (a) and (b) struggle with either precise structural preservation or faithful content generation, our DVD (c) can yield flat document images with precise yet faithful content through a mapping generation paradigm.
Document dewarping has evolved through two distinct periods. The early model-based methods (Kanungo et al., 1993; Cao et al., 2003a; Liang et al., 2008) basically follow a ”reconstruct-then-dewarp” paradigm, which is typically limited by hardware dependence or surface assumptions. Subsequent data-driven methods (Ma et al., 2018; Li et al., 2019; Das et al., 2019) have shifted toward a mapping regression paradigm to model the condition probability. Specifically, given a large-scale photographic document dataset, these methods predominantly train neural networks to directly regress mapping (e.g., backward mapping, forward mapping) for dewarping. Accordingly, they also transfer the research attention to high-fidelity training data (Zhang et al., 2024; Verhoeven et al., 2023) and better network architecture for feature extraction (Li et al., 2023b; Feng et al., 2022).
Albeit accomplishing notable advancements, the contemporary mapping regression paradigm is burdened by its intrinsic discriminative nature, which lacks the capacity to explicitly capture the underlying data distribution, resulting in imprecise structural preservation (See Fig. 1 (a)). Recent diffusion models have demonstrated the viability of employing a generative paradigm to solve discriminative tasks (Ravishankar et al., 2024; Nam et al., 2024; Luo et al., 2024). By learning a progressive denoising process to generate samples conforming to the training data distribution, these models introduce a generative task modeling to learn more comprehensive distributions. Most recently, MLLMs (e.g., Gemini 2.5, GPT-4o) have exhibited remarkable natural image generation capabilities (Yan et al., 2025), but failed to preserve document structures in document dewarping (e.g., Fig. 1 (b)). We attribute this to the fact that it’s very difficult to directly employ an image translation paradigm to obtain faithful control in highly complex document images. In light of these insights, we propose one intriguing question: Can generative models be effective for document dewarping?
To answer this question, we present DvD, the first effort to unleash a generative dewarping model built on a coordinates-based denoising diffusion framework. Instead of typical pixel-level denoising, we introduce a coordinate-level denoising process, where DvD learns how to progressively generate a series of latent variables to characterize the mapping for deformation rectification. We argue that such a mapping generation paradigm not only explicitly fosters deformation-aware modeling but also avoids the difficulty of high-resolution image generation. To further enhance the structural preservation, DvD also incorporates a time-variant condition refinement mechanism to leverage intermediate dewarping results. Diverging from typical diffusion models guided by a time-fixed condition, our proposed time-variant mechanism enables intermediate-aware dynamic guidance in the denoising generation process.
To make the evaluation of dewarping models fairly, we collect most of publicly available document dewarping benchmarks as shown in Table 1. We find that these benchmarks typically suffer from three critical limitations that impede a comprehensive evaluation of dewarping models, including restricted coverage of real-world scenarios, small scale of dataset size, and deficient annotation of domains. These limitations might have led to evaluation ambiguity, restricting the model’s application in real-world scenarios. To this end, we construct a fine-grained and large-scale benchmark AnyPhotoDoc6300, which contains 6,300 real-world photographic image document pairs, rigorously organized across three distinct domains, enabling fine-grained evaluation of dewarping models. In addition to the benchmark, we extend the evaluation metrics. We find that existing methods commonly utilize off-the-shelf OCR engines to measure the Edit Distance (ED) and Character Error Rate (CER) as OCR metrics (Das et al., 2019). We identify that there is still no exploration about whether the dewarped document can attain equivalent text readability to its flat counterpart for MLLMs. To fill the blank, we propose two MLLM-based OCR metrics (MMCER and MMED) in this work.
In summary, our main contributions are four-fold:
  • Pioneering a paradigm shift, we present DvD, the first generative model to tackle document dewarping via a coordinates-based diffusion framework. Unlike typical pixel-level denoising to generate dewarped images directly, we operate coordinate-level denoising to generate coordinates mappings for dewarping.
  • We introduce a time-variant condition refinement mechanism that dynamically leverages intermediate dewarping results as guidance to enhance the preservation of document structures.
  • To offer a comprehensive evaluation of document dewarping models, we construct a fine-grained benchmark dataset AnyPhotoDoc 6300, which contains 6,300 real-world photographic image document pairs, rigorously organized across three distinct domains.
  • Extensive experiments demonstrate state-of-the-art dewarping performance with acceptable computational efficiency.
Table 1. Comparison of document dewarping benchmarks. ✗ symbolizes that this benchmark doesn’t explicitly distinguish this domain. Scenes: Multiple scenarios (Mul), Real (R), Synthetic (S). Domains: Layouts Category (L), Environment Lighting (E), Capture Angles (A).
Dataset Scenes Images Domains
L E A
DocUNet (Ma et al., 2018) Mul-R $130\times 2$
DIR300 (Feng et al., 2022) Mul-R $300\times 2$
Inv3DReal (Hertlein et al., 2023) Invoice-R $360\times 2$ 1 3 1
UVDoc (Verhoeven et al., 2023) Mul-S $50\times 2$
DocReal (Yu et al., 2024) Chinese-R $200\times 2$
Our AnyPhotoDoc6300 Mul-R $\textbf{6300}\times 2$ 8 3 2
$\bullet$ Noted that we don’t list WarpDoc (Xue et al., 2022) and SP (Li et al., 2023a)
due to open-source incompleteness.
2. Related work
2.1. Early Model-based Methods
Early model-based methods basically adhere to a two-step ”reconstruct-then-dewarp” paradigm. At the first step, for surface reconstruction, some methods utilize specialized hardware such as structured-light devices (Brown and Seales, 2001; Meng et al., 2014), laser scanners (Zhang et al., 2008), multi-view camera systems (Ulges et al., 2004; Koo et al., 2009; You et al., 2017a), and proximal light sources (Wada et al., 1997). Another methods bypass the hardware dependence by leveraging 2D visual cues (e.g., text lines (Tian and Narasimhan, 2011), local text orientation (Meng et al., 2018), text blocks (Kim et al., 2015), etc.) to estimate 3D geometry with parametric assumptions. In the second step, hand-crafted transformations for dewarping the surface to the plane are undertaken according to the reconstructed surface. For parameterized surfaces, typical transformations involve Generalized Cylinders Surface (GCS) (Cao et al., 2003b; Koo et al., 2009), developable surfaces (Liang et al., 2008), generalized ruled surfaces (Tsoi and Brown, 2007), and Non-Uniform Rational B-Splines (NURBS) (Tan et al., 2005). For non-parameterized surfaces, techniques such as planar strips (Meng et al., 2015), stiff mass-spring systems (Zhang et al., 2008), conformal mapping (You et al., 2017a), and sparse correspondence (Meng et al., 2018) are harnessed to tackle the mesh representation.
Figure 2. Framework of the proposed DvD. Given a warped document as input, DvD generates latent variables $m$ to characterize the mapping for deformation rectification. We leverage a compound condition $c_{t}$: the features for raw document images $f_{d}$, document foreground mask $f_{m}$, text-lines $f_{l}$, and the time-variant condition of the intermediate result $r_{t}$, respectively. We visualize the reverse denoise generation process in the right pink region.
While effective in limited scenarios, these methods exhibited poor generalization to real-world warping patterns due to confined assumptions. Meanwhile, demanding hardware-dependent solutions also deviate from the prevalent habit of using mobile phones to digitize documents.
2.2. Data-driven Methods
Data-driven methods bring a shift toward a regression-based paradigm, significantly lessening reliance on both assumptions and hardware. Ma et al. (2018) pioneer to employ deep neural networks to solve document dewarping by framing it as a mapping regression task. Subsequent DewarpNet (Das et al., 2019) reformulates the traditional ”reconstruct-then-dewarp” paradigm via two regression networks while collecting a high-quality synthetic dataset Doc3D using rendering software. CREASE (Markovitz et al., 2020) explores multi-modal warped document representations to strengthen the dewarping performance. DocProj (Li et al., 2019) and PW-DewarpNet (Das et al., 2021) opt to dewarp documents in a patch-wise manner, benefiting the results at local details. DispFlow (Xie et al., 2020) and DDCP (Xie et al., 2021) investigate the effectiveness of mapping and sparse control points as deformation representations, respectively. DocTr (Feng et al., 2021) replaces convolutional network with vision transformer, achieving significant performance boosts. RDGR (Jiang et al., 2022), DocGeoNet (Feng et al., 2022), and FTA (Li et al., 2023b) leverage text-line features to keep textual content alignment. PaperEdge (Ma et al., 2022), Marior (Zhang et al., 2022), and DocReal (Yu et al., 2024) devise two-stage networks, which first coarsely dewarp the mask of document, then refine details. FDRNet (Xue et al., 2022) introduces frequency-domain insight, designing an image-level loss based on high-frequency textures extracted from the Fourier transformation. UVDoc (Verhoeven et al., 2023) exploits a novel data annotation pipeline through optical invisibility of ultraviolet ink to acquire pseudo-real training data. LA-Doc (Li et al., 2023a) constructs a synthetic dataset with large-scale deformation patterns via physical mass-spring systems and proposes a layout-aware dewarping network. Inv3D and DocMatcher (Hertlein et al., 2025, 2023) present a template-based document dewarping using auxiliary template information. These methods concentrate on curating training datasets and devising regression networks, whereas their intrinsic discriminative nature is unexplored, which constrains the dewarping performance.
3. Methodology
3.1. Document Dewarping Framework
As illustrated in Fig. 3, given a warped photographic document image $I_{w}\in\mathbb{R}^{H\times W\times 3}$ as input, document dewarping aims to restore the flat document texture $I_{0}\in\mathbb{R}^{H\times W\times 3}$. In this work, we firstly propose a dewarping model to predict a small-scale coordinates mapping $m_{0}$, saving computational cost, which is then up-sampled to a normal backward mapping $M_{0}\in\mathbb{R}^{H\times W\times 2}$ (each value in $M_{0}$ represents the corresponding 2D coordinates in the input warped image $I_{w}$ as shown in Equ. 1), Next, we employ a dense spatial transformation $T_{m}$ to calculate the pixel values in $I_{w}$ according to the coordinates from $M_{0}$, ultimately obtaining $I_{0}$ as shown in Equ. 1. We formalize the backward mapping procedure as
(1) $\begin{split}(i,j)&=M_{0}(u,v),\\ I_{0}(u,v)&=T_{m}(I_{w}(i,j)),\end{split}$
where $(i,j)$ and $(u,v)$ represent the 2D spatial coordinates of $I_{w}$ and $I_{0}$, respectively. In the following, we will describe our dewarping model in details.
Figure 3. General document dewarping pipeline. We propose a dewarping model to predict a backward mapping for deformation rectification, consisting of two coordinates channels.
3.2. Coordinates-Based Diffusion Model
Latent Diffusion Model in Coordinate Space. Directly generating $M_{0}$ explicit foster deformation-aware modeling to achieve better structural preservation, however both $M_{0}$ and $I_{0}$ actually contain equally high-resolution (e.g., 2000×3000), leading to high computational complexity. Inspired by the latent diffusion model (Rombach et al., 2021), we implement the diffusion and denoising processes only in a smaller coordinate space (i.e., $64\times 64$), significantly reducing the training and inference costs. Figure 2 shows the overall framework of our proposed DvD model. Concretely, guided by a compound condition $c_{t}$, our DvD progressively generates a series of latent variables $m$ from a random Gaussian distribution. This paradigm relies on a tailored conditional denoising diffusion probabilistic framework, which typically contains both forward and reverse processes (Ho et al., 2020; Song and Ermon, 2019). We define the forward diffusion process for mapping as the Gaussian transition, s.t. $q(m_{t}|m_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}m_{t-1},\beta_{t}\mathbf{I})$, where ${\beta_{t}}$ is a predefined variance schedule. The resulting latent variable $m_{t}$ can be formulated as:
(2) $m_{t}=\sqrt{\bar{\alpha}_{t}}m_{0}+\sqrt{1-\bar{\alpha}_{t}}z,\quad z\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\left(1-\beta_{i}\right)$, and $m_{0}$ is the ground-truth mapping. Afterward, following Nichol and Dhariwal (2021), we train a neural network $\mathcal{\epsilon}_{\theta}(\cdot)$ for the reverse denoising process, during which the initial latent variable ${m}_{T}$ is iteratively denoised following the sequence ${m}_{T-1},{m}_{T-2},\ldots,{m}_{0}$, as follows:
(3) $\begin{split}m_{t-1}=&\sqrt{\bar{\alpha}_{t-1}}\mathcal{\epsilon}_{\theta}(m_{t},t,c_{t})+\\ &\frac{\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}}{\sqrt{1-\bar{\alpha}_{t}}}\Bigl{(}m_{t}-\sqrt{\bar{\alpha}_{t}}\mathcal{\epsilon}_{\theta}(m_{t},t,c_{t})\Bigr{)}+\sigma_{t}z,\end{split}$
where $\mathcal{\epsilon}_{\theta}(m_{t},t,c_{t})$ directly predicts the denoised mapping $\hat{m}_{0,t}$.
Compound Conditions $c_{t}$. Following previous works (Feng et al., 2022; Li et al., 2023b; Jiang et al., 2022), our DvD utilizes pre-trained multiple feature extractors (MFE) to enhance document visual perception and composes these features into a compound condition for the diffusion model, denoted as $c_{t}=\{f_{d},f_{m},f_{l},r_{t}\}$, where $f_{d},f_{m},f_{l}$ represent the features for raw document images, document foreground mask and text-lines, respectively. Note that $f_{d},f_{m},f_{l}$ are time-fixed conditions, while $r_{t}$ represents a time-variant condition that enables dynamic guidance in the denoising generation process.
Time-variant Condition Refinement (TVCR) Mechanism. As illustrated in the right region of Fig. 2, the intermediate dewarping results reveal the visual gap from intermediate denoising states to the ideal dewarped result. To harness this information for enhanced the preservation of document structures, we introduce a time-variant condition refinement mechanism within the reverse diffusion process to ensure increasingly precise document structure as the process evolves. Specifically, we incorporate a time-variant condition $r_{t}=\{m_{0|t},f_{0|t}\}$ for iteratively updating $r_{t}$, where $m_{0|t}$ means the predicted latent variables mapping in each time-step, and $f_{0|t}$ indicates dewarped document features $f_{d}$ using $m_{0|t}$ by Equ. 1. Unlike the time-fixed condition applied in vanilla diffusion models, the proposed time-variant condition reflects the intermediate variable $m_{0|t}$ and local structure deformation $f_{0|t}$ in each diffusion step, which facilitates the model to dynamically approach a tighter evidence lower bound (ELBO) (Rezende et al., 2014) for maximizing the conditional likelihood $\log p(m_{0}|\{c_{t}\})$.
Figure 4. A sample data array visualization across 3 distinct domain spectra (A, B, C) in our AnyPhotoDoc6300 Benchmark. We aim to evaluate dewarping capability of the model to different warping patterns (D) under well-distinguished domain combinations.
Figure 5. Quantitative dewarping performance comparisons on the AnyPhotoDoc6600 benchmark dataset. Under the fixed ”Two-column Paper” Layout Pattern, we provided 18 dimensions of evaluation for AD, MS-SSIM, and MMCER. More evaluation results can be found on the Figures-only pages after the reference.
3.3. Network Architecture
In this section, we discuss the design of the network architecture in Fig. 2, including the multiple feature extractors (MFE) and diffusion decoder. In MFE, we employ three parallel networks to extract features $f_{d},f_{m},f_{l}$ for raw document image, document foreground, and text lines, respectively. For document image features $f_{d}$, we utilize the first three layers of VGG16 (Simonyan and Zisserman, 2014) to extract features. For document foreground features $f_{m}$, we concatenate the last six layers of U2Net (Qin et al., 2020) as a foreground segmentation network to extract those features associated with foreground. For text lines features $f_{l}$, we use the last decoder layer of the UNet (Ronneberger et al., 2015), which is pre-trained for text-line segmentation. All extracted features are uniformly downsampled to $64\times 64$ resolution compatible with the coordinate space. Furthermore, the diffusion decoder comprises 12 condition embedding blocks (CEB) and 6 fusion generation blocks (FGB). In each condition embedding block, we extend the architecture of diffusion transformers (DiT) (Peebles and Xie, 2023) to implement both time embeddings and cross-attention conditioning. The three features $f_{d},f_{m},f_{l}$ from the MFE and time-variant conditions ($r_{t}$) are jointly conditioned with the noisy latent variable $m_{t}$ through four parallel cross-attention computations. In the FGB, the results of the cross-attention computations are concatenated and fed into a six-layer FGB that contains self-attention and feed-forward layers. Subsequently, we apply three linear layers to obtain the denoised mapping $m_{0}$. Next, we upsample $m_{0}$ to obtain the high-resolution mapping $M_{0}$, which is used to produce the final flat document image $I_{0}$ via Equ. 1. In Appendix B, we provide more detailed architectures.
3.4. Training and Sampling
DvD Training Algorithm. Vanilla diffusion model randomly samples timesteps from a uniform distribution during the training phase (Ho et al., 2020), which mismatch the sequential acquisition of time-variant conditions. To solve this issue, we extend the vanilla diffusion model by integrating our TVCR mechanism, as detailed in Algorithm LABEL:alg:training, where we implement a certain number of sampling based on current timestep, thereby obtaining the intermediate dewarping latent variables $m_{0|t}$ and the corresponding intermediate dewarped features $f_{0|t}$.
Model Optimization. During the training phase, we freeze the pre-trained weights of the foreground and text-line segmentation network from DocGeoNet (Feng et al., 2022), including U2Net and UNet. However, for the VGG dedicated to document feature extraction raw document feature extraction, we jointly optimize it with the subsequent diffusion decoder. To optimize our DvD, instead of predicting noise in (Ho et al., 2020), we follow Luo et al. (2024) and Nam et al. (2024) to predict the generated object itself. Thus, the loss function is given by:
(4) $\begin{split}\mathcal{L}=\mathbb{E}_{m_{0}\sim q(m_{0}|c_{t}),z\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\left\|m_{0}-\mathcal{\epsilon}_{\theta}\left(m_{t},t,c_{t}\right)\right\|^{2}\right],\end{split}$
Stochastic Sampling Property. As shown in Equ. 3, the reverse process of DDIM (Nichol and Dhariwal, 2021) introduces $\sigma$ to inject stochasticity into the sampling trajectory. To account for this property and enhance generation stability, we implement a dual-hypothesis strategy that simultaneously generates two candidate mappings. Afterward, we calculate the mean of the two mappings as a final result. We provide further training details in Appendix A.
             (Input)                   (Das et al., 2019)             (Ma et al., 2022)             (Feng et al., 2022)         (Verhoeven et al., 2023)         (Li et al., 2023b)             (March,25th,2025)                  (Our)
Figure 6. Qualitative comparisons on the AnyPhotoDoc6300 benchmark. We highlight some obvious content edges with red dotted lines. More visual comparisons can be found on the Figures-only pages after the reference.
Table 2. Quantitative dewarping performance comparisons on the DocUNet benchmark dataset. Bold indicates the best, underline indicates second-best. The last column shows the network size by the number of parameters (millions).
Method Venue Training Dataset MS-SSIM $\uparrow$ LD $\downarrow$ AD $\downarrow$ CER $\downarrow$ ED $\downarrow$ MMCER $\downarrow$ MMED $\downarrow$ Para.
Warped Document - - 0.246 20.51 1.026 0.595 1819.16 0.576 700.96 -
Training under Non-uniform or Proprietary Dataset
DispFlow (Xie et al., 2020) DAS’20 DIWF 0.431 7.64 0.411 0.446 1322.94 0.887 1339.43 23.6M
DDCP (Xie et al., 2021) ICDAR’21 DDCP 0.474 8.92 0.459 0.458 1335.30 0.655 762.28 13.3M
PaperEdge (Ma et al., 2022) SIGGRAPH’22 Doc3D+DIW 0.472 8.01 0.385 0.407 1038.55 0.198 530.27 36.6M
UVDoc (Verhoeven et al., 2023) SIGGRAPHA’23 Doc3D+UVDoc 0.544 6.83 0.315 0.384 1026.91 0.402 615.88 8M
LADoc (Li et al., 2023a) TOG’23 Doc3D+SP 0.523 7.24 0.310 0.395 956.27 0.242 518.74 -
DocReal (Yu et al., 2024) WACV’24 Doc3D+AugDoc3D 0.502 7.00 0.284 0.394 1032.17 0.336 547.05 -
Training under Uniform Doc3D Dataset
DewarpNet (Das et al., 2019) ICCV’19 Doc3D 0.472 8.41 0.412 0.441 1158.66 0.533 734.84 86.9M
DocTr (Feng et al., 2021) MM’21 Doc3D 0.511 7.77 0.365 0.403 1093.66 0.432 615.38 26.9M
RDGR (Jiang et al., 2022) CVPR’22 Doc3D 0.496 8.53 0.453 0.372 994.01 0.403 534.97 -
Marior (Zhang et al., 2022) MM’22 Doc3D 0.448 8.42 0.470 0.421 1131.48 0.232 510.11 -
DocGeoNet (Feng et al., 2022) ECCV’22 Doc3D 0.504 7.69 0.378 0.367 993.08 0.376 627.11 24.8M
FTA (Li et al., 2023b) ICCV’23 Doc3D 0.494 8.87 0.391 0.403 1093.63 0.355 544.11 45.2M
DvD - Doc3D 0.549 6.61 0.279 0.366 928.94 0.215 515.97 151.25M
4. Experiments
4.1. AnyPhotoDoc6300 Benchmark
Despite significant advances in document dewarping, the development of corresponding benchmarks lag behind. We summarized current benchmarks in Tab. 1, and we can find most of datasets suffer from restricted coverage of scenarios, small size of dataset, and deficient annotations of domains, impeding a comprehensive evaluation of dewarping models. To this end, we build a new large-scale benchmark AnyPhotoDoc6300, containing 6,300 real-world photographic document pairs. In our AnyPhotoDoc6300, we provide three distinct domain annotations to enable a more fine-grained quantitative evaluation, including layout category (LC), environment lighting (EL), and capture angles (CA). In addition, we aim to evaluate the dewarping capability of the model on three typical warping patterns (i.e., curves, folds, and crumples) under any combination of three types of given domains. Fig. 4 visualizes a sample array of ”Complex layout paper” (one of the eight layout categories). By meticulously specifying three environmental lighting and two capture angles, we can form any domain combination for the same document content. Consequently, we can obtain a fine-grained performance evaluation like Fig. 5 to discover the underlying issues for the current dewarping model. In Appendix E, we provide more details about our AnyPhotoDoc6300 benchmark, including collection settings and more visualization from different layout categories. We put some examples of AnyPhotoDoc6300 benchmark in supplementary materials and the completed benchmark will be publicly available after the paper review.
4.2. Metrics
Feature Alignment metrics. Following most of previous models (Ma et al., 2018; Das et al., 2019; Ma et al., 2022), we adopt the MS-SSIM (Multi-scale Structural Similarity) (Wang et al., 2003), LD (Local Distortion) (You et al., 2017b), and AD (Aligned Distortion) (Ma et al., 2022) to evaluate differences between dewarped and flat ground truth. Among them, MS-SSIM focuses on perceiving similarity on luminance, contrast, and structural information. While LD and AD focus on measuring the variations of SIFT flow (Liu et al., 2010).
OCR metrics. To evaluate the OCR readability improvement enabled by document dewarping, measuring recognition discrepancy between dewarped and flat documents has become the de facto standard in contemporary document dewarping models. Concretely, an off-the-shelf OCR engine is applied to recognize two text sequences from dewarped and flat documents, respectively. Then, ED (Edit Distance) and CER (Character Error Rate)  (Levenshtein et al., 1966) are harnessed to quantify the degree of deviation between two sequences. Pioneering a metric supplement, we further extend current OCR metrics in document dewarping by replacing previous OCR engines with MLLMs. Witnessing the superior progress of MLLMs in OCR capabilities recently (Wei et al., 2024; Karmanov et al., 2025; Nassar et al., 2025), we identify that there is still no exploration about whether the dewarped document can attain equivalent readability to its flat counterpart for MLLMs. To fill the blank, we pioneer MLLM-based OCR metrics (i.e., MMCER, MMED) to serve as a specialized supplement for prevalent ED and CER. Specifically, given the prompt of ”OCR the plain text” as a fixed instruction, we employ open-source MLLM QWen2.5-VL 7B (Bai et al., 2025) to recognize all characters in both dewarped and flat documents for ED and CER calculation.
4.3. Qualitative and Quantitative Comparison
Qualitative Comparisons. Dewarped document results on the AnyPhotoDoc6300 and DocUNet benchmark are shown in Fig. 6 and Figure-only pages. In Fig. 6, we select the five most recent open‑source dewarping models as well as GPT‑4o’s native generation model. On figure‑only pages, we additionally compare against publicly released inference results from non‑open‑source methods. Basically, our DvD achieves precise structure preservation in both local and overall document content. In contrast, GPT‑4o tends to produce unfaithful results that look visually clean but whose content is often chaotic. We argue this is because the image translation paradigm adopted by GPT-4o lacks explicit deformation awareness.
Quantitative Comparisons. We compare the performance of our DvD model with previous state-of-the-art on three benchmarks, including DocUNet (Ma et al., 2018), DIR300 (Feng et al., 2022), and the proposed AnyPhotoDoc6300. The quantitative results on the DocUNet benchmark are shown in Tab. 2. For DIR300, we have placed the results in Appendix C. The ”Warped Document” in the first row means that we simply feed the raw input image for evaluation, therefore, most of the metrics perform the worst. Since our paper focuses on novel model designs rather than contributing high-quality data, we only train DvD on the currently most widely used Doc3D (Das et al., 2019) dataset. For a fair comparison, we divide the current methods into two branches according to their training data. It can be seen that even though we only used the Doc3D dataset, our method still achieved the best performance on the majority of metrics. In the upper first branch, DvD can achieve a slight overtaking, while in the lower second branch, DvD can achieve a significant superiority under a uniform Doc3D dataset. To be noted, compared with the first row, our experiments on MMCER and MMED reveal counterintuitive performance degradation in earlier dewarping models (e.g., DispFlow, DewarpNet), which we attribute to the fact that MLLMs are sensitive to resolution reduction and artifact (Li et al., 2024; Feng et al., 2024). On the other hand, compared with the first row, our DvD significantly reduces MMCER and MMED, and greatly improves the perception of photographed documents for MLLMs.
Fine-grained Quantitative Comparisons. Fig. 5 illustrates AD, MS‑SSIM, and MMCER as three representative metrics via the Any-PhotoDoc6300 benchmark. More metrics in different domains are also exhibited in the Figures-only pages. Our DvD comprehensively achieves superior performance across various domains, including layout, lighting, and angles. Moreover, for the first time, we can also unveil performance of dewarping models at a fine-grained level. For instance, In Fig. 5 (I), existing methods generally suffer from poor AD performance on the warping pattern of the curves under moderate or warm lighting. In Fig. 5 (II), we observe a notable MS-SSIM drop for dark lighting documents with folds and crumples. In Fig. 5 (III), warm lighting causes the most pronounced decline in OCR metrics, especially on documents with curves. By pinpointing these issues unconventionally, we expect to provide the research community deeper insights into dewarping model behaviors and dataset curation, for driving further performance improvements.
4.4. Ablation Studies
Effectiveness on different Dewarping Paradigms. To verify the superiority of the proposed paradigm with the regression-based paradigm, we specially train another network of DvD by directly regressing the mapping. Then we can fairly compare different learning paradigms under the same network structure. As demonstrated in Tab. 3, The DvD baseline model trained using the regression-based paradigm has led to a general decline in performance, which emphasizes the effectiveness of our mapping generation paradigm for obtaining a more precise structure preservation.
Component Analysis of Compound Conditions. Tab. 4 demonstrates three types of compound condition $c_{t}$ configurations. We cam see that only using raw document feature $f_{d}$ does not obtain satisfied performance, which is then improved by adding document foreground $f_{m}$ and text-lines $f_{l}$. Finally, adding time-variant condition $r_{t}$ further boosts the performance obviously on all metrics. All of these verify the complementarity of those compound conditions.
Table 3. Ablation study on different learning paradigms.
Learning paradigms MS-SSIM $\uparrow$ LD $\downarrow$ AD $\downarrow$ CER $\downarrow$ ED $\downarrow$
DvD (Mapping Regression) 0.487 7.85 0.482 0.410 1084.67
DvD (Mapping Generation) 0.549 6.61 0.279 0.366 928.94
Table 4. Ablation study for different conditions components
Conditions $c_{t}$ Experimental Results
[…]
$\checkmark$ 0.409 8.18 0.332 0.427 1184.52
$\checkmark$ $\checkmark$ 0.519 7.12 0.323 0.408 1046.71
$\checkmark$ $\checkmark$ $\checkmark$ 0.549 6.61 0.279 0.366 928.94
Table 5. Ablations on diffusion step, performance, and time consumption.
[…]
1 0.420 9.86 0.501 0.875 1952.54 0.21
3 0.549 6.61 0.279 0.366 928.94 0.59
5 0.537 6.60 0.281 0.372 956.46 1.06
50 0.475 7.56 0.467 0.441 1261.43 10.32
Computational Efficiency Analysis. Tab. 5 illustrates a trade-off between performance and time consumption, driven by different diffusion denoising steps. As the number of steps increases from 1 to 3, the model’s performance improves substantially. However, beyond 3 steps, performance largely plateaus or even slightly declines. At the same time, increasing the step count incurs a linearly growing time cost. In our experiments, we set 3 as the optimal step setting for a trade off between performance and time consumption.
5. Conclusion
This paper unleashes a novel mapping generation paradigm for the document dewarping task by reformulating it as a coordinate-based denoising diffusion framework. To the best of our knowledge, this is the first attempt to explore the viability of dewarping a document using the diffusion model, where we tailor a coordinate‑level denoising strategy and a time-variant condition refinement (TVCR) mechanism, enabling precise preservation of document structures. To foster a fine-grained evaluation of dewarping models, we also build a new photographic document dewarping benchmark AnyPhotoDoc 6300, which is large-scale size, covers multiple scenarios and provides detailed domain annotations. Findings and insights from our experiments are poised to substantially advance photographic document processing and further impact a broad spectrum of graphics applications.